# COREval: A Comprehensive and Objective Benchmark for Evaluating the Remote Sensing Capabilities of Large Vision-Language Models

### v1

@@ 5 Conclusion
In this work, we introduce COREval, the first benchmark for comprehensively and objectively evaluating the remote sensing capabilities of VLMs. Through COREval, we assessed multiple mainstream general-domain VLMs and RSVLMs, across hierarchical capability dimensions. Our multi-dimensional analysis reveals the strengths and limitations of these models, providing insights to enhance their effectiveness in remote sensing applications. This work not only establishes a valuable resource for benchmarking VLMs in remote sensing but also offers strategic guidance for future advancements in the field.
Supplementary Material
6 Evaluation Dimension
To thoroughly evaluate the perception and reasoning capabilities of VLMs, COREval is structured with 6 L-2 dimensions and 22 L-3 leaf tasks. In this section, we provide a detailed definition of each L-3 task, along with a collection of illustrative examples. Following this, we present a detailed statistical summary of COREval.
6.1 Definition of Each L-3 Task
Image-level Comprehension. This evaluation dimension contains 5 L-3 tasks, focusing on a broad understanding of image-level content and field-specific relevance. In Fig. 3, we present some visualization examples for this dimension. Notably, the titles in black indicate that the L-3 tasks are constructed by the label-driven method, titles in red signify tasks developed through the foundation model-driven method, and titles in purple denote tasks created through human-GPT-4 collaboration.
@@ 6.1 Definition of Each L-3 Task
Single-instance Identification. This dimension emphasizes detailed object recognition and localization without reliance on prior knowledge or fixed categories, including 7 L-3 tasks. In Fig. 4, we present some visualization examples for this dimension.
@@
Cross-instance Discernment. This dimension emphasizes the comparisons or changes among multiple objects of interest, which provides insights into contextual dependencies, spatial dynamics and temporal changes within an image. Three L-3 tasks are included and Fig. 5 presents some visualization examples for each task.
@@
Attribute Reasoning. In this dimension, VLMs are required to reason about the overall characteristics of RSIs or the specific attributes of the instances. This dimension includes 2 Level-3 tasks, with visualization examples for each task shown in Fig. 6.
@@
Assessment Reasoning. In this dimension, VLMs are required to reason about estimated indicators of societal development and environmental conditions depicted in the RSI. This dimension includes 2 Level-3 tasks, with visualization examples for each task shown in Fig. 7.
@@
    Resource Assessment: This task evaluates a VLM’s capability to deduce the estimated resource quantity present in a given area, such as the population sizes.
@@
Common Sense Reasoning. In this dimension, VLMs are required to reason about real-world scenarios based on the extensive common sense knowledge embedded in LLM. This dimension includes 3 Level-3 tasks, with visualization examples for each task shown in Fig. 8.
@@
6.2 Statistics of COREval
All RSIs in COREval are sourced from a variety of satellites, products, and platforms. To ensure a balanced evaluation, we targeted an even distribution among problems related to different capabilities during data collection, with a minimum of 100 samples per task. The data sources and the number of problems are listed in Tab. 5.
7 Details on COREval Construction
To increase data diversity, most RSIs in COREval are sourced from 50 randomly selected cities across six continents (excluding Antarctica), based on the top 1,000 global cities identified by Oxford Economics’ Global Cities Index, minimizing potential biases in the data. The details of data coverage are listed in Tab. 6
     Level-1 Level-2 Level-3 Data Source Question Type # Problem
@@ 7 Details on COREval Construction
Sentinel-1
2-option MCQ
4-option MCQ
700
MR Natural Earth 4-option MCQ 157
SC Google Earth 4-option MCQ 2000
IC Google Earth 4-option MCQ 200
SII LR Google Earth 4-option MCQ 100
OC Google Earth 4-option MCQ 200
OL Google Earth 4-option MCQ 200
OP Google Earth 4-option MCQ 200
AR SWISSIMAGE [^4] https://www.swisstopo.admin.ch/de/en/orthoimage-swissimage-10 4-option MCQ 144
VG Google Earth Free-form QA 200
HD Google Earth 4-option MCQ 200
CID AC SWISSIMAGE 2-option MCQ 205
SR SWISSIMAGE 4-option MCQ 157
CD Google Earth 4-option MCQ 200
Reasoning AttR TP Sentinel-2 4-option MCQ 200
@@
Google Earth
CNBH [42]
4-option MCQ 100
@@
Google Earth
ODIAC [30]
4-option MCQ 200
@@
4-option MCQ 200
CSR DD Diffusion Model 4-option MCQ 200
GD Google Earth 4-option MCQ 200
SI Google Earth 4-option MCQ 100
       Total 6,263
Table 5: The detailed statistics for COREval.
@@
For Scene Classification, we begin by defining a set of categories for RSIs to guide data collection. OpenStreetMap (OSM) represents physical features on the ground (e.g., roads or buildings) using tags (i.e., key-value pairs), with each tag describing a geographic attribute of the feature. For example, an aerodrome, airport or airfield can be tagged as $aeroway=aerodrome$. Take the common categories in existing relevant datasets [43, 5] as a reference, we follow the OSM’s tagging system to establish a category set, as shown in Tab. 7, with each predefined category can be described by one or more key-value pairs. We further select 10 of the 50 chosen cities as the primary data collection areas to capture more intra-class variations, including Sapporo, Mecca, Busan, Shanghai, Hamburg, Amsterdam, Phoenix, Cairo, Santiago and Melbourne. As shown in Fig. 9, we obtain polygons representing the locations of interest based on the predefined key-value pairs for each category. Using these polygons, we then download the corresponding RSIs from Google Earth. To ensure high-quality data, we design an automatic filtering pipeline to exclude the low-confidence RSIs resulting from the inaccuracy in OSM. Following [2], we finetune two versions of classification models based on the 30-category AID dataset and 45-category NR dataset, both of which include the 20 categories we defined. After cropping the RSIs we collect into $512\times 512$ tiles, we input them into the two models and take the intersection of the outputs with over 99% confidence from both models as the desired RSIs. Finally, human annotators review the results to further ensure the accuracy of all categories. We ultimately collect 2,000 high-confidence RSIs across 20 categories for this task.
For Image Quality, we first define 4 common types of noise that typically interfere with RSIs: Gaussian noise, impulse noise, stripe noise and deadline noise. We randomly select 200 samples in Scene Classification evenly as the original high-quality RSIs, and manually add each type of noise to these samples, creating 50 samples for each noise type, resulting in 200 multiple-choice questions (MCQs) for identifying the type of noise present in an RSI. Moreover, we concatenate each noisy RSI with its corresponding original RSI side by side, forming 200 two-option MCQs to determine which version has better quality. We further collect cloudy RSIs from the Landsat-8 satellite, which has a spatial resolution of 30 meter/pixel, across the 10 cities. By selecting tiles with 60% cloud coverage and cropping the corresponding areas into $512\times 512$ patches, we combine Bands 4, 3 and 2 to obtain the natural color images, which results in 100 cloudy RSIs. We then concatenate three RSIs from Time Property and one cloudy RSI into a $2\times 2$ grid, and generate 100 MCQs for identifying the cloudy RSI. To comprehensively assess the VLM’s capability to identify and compare the quality of RSIs, we concatenate the original RSI, its corresponding noisy version, the cloudy RSI and the low-resolution RSI (obtained by randomly cropping the original RSI and resizing it to the original size) into a $2\times 2$ grid, forming 200 MCQs to identify the best quality among the four options.
For Image Modality, we include four modalities of RSIs: RGB, SAR, false color and nighttime light. We still select the RSIs from Scene classification as the RGB group, and combine Bands 5, 4 and 3 of the cloudy RSIs from Landsat-8 as the false color modality. With the assistance of the Sentinel-1 satellite with a spatial resolution of 10 meter/pixel, we download and crop SAR RSIs across the 10 cities. As for nighttime light images, we manually choose RSIs from the SDGSAT-1 satellite. We then randomly select 50 RSIs per modality to form 200 MCQs for identifying the modality of each RSI.

### v2

@@ 5 Conclusion
In this work, we introduce CHOICE, the first benchmark for systematically and objectively evaluating the remote sensing capabilities of VLMs. Through CHOICE, we assessed 24 mainstream general-domain VLMs and RSVLMs across hierarchical capability dimensions. Our multi-dimensional analysis reveals the strengths and limitations of these VLMs, providing insights to enhance their effectiveness in remote sensing applications. This work not only establishes a valuable resource for benchmarking VLMs in remote sensing but also offers strategic guidance for future advancements in the field.
Appendix A Evaluation Dimension
To thoroughly evaluate the perception and reasoning capabilities of VLMs, CHOICE is structured with 6 L-2 dimensions and 23 L-3 leaf tasks. In this section, we provide a detailed definition of each L-3 task, along with a collection of illustrative examples. Following this, we present a detailed statistical summary of CHOICE.
A.1 Definition of Each L-3 Task
Image-level Comprehension. This evaluation dimension contains 5 L-3 tasks, focusing on a broad understanding of image-level content and field-specific relevance. In Figure 4, we present some visualization examples for this dimension. Notably, the titles in black indicate that the L-3 tasks are constructed by the label-driven method, titles in red signify tasks developed through the foundation model-driven method, and titles in purple denote tasks created through human-GPT-4 collaboration.
@@ A.1 Definition of Each L-3 Task
Single-instance Identification. This dimension emphasizes detailed object recognition and localization without reliance on prior knowledge or fixed categories, including 7 L-3 tasks. In Figure 5, we present some visualization examples for this dimension.
@@
Cross-instance Discernment. This dimension emphasizes the comparisons or changes among multiple objects of interest, which provides insights into contextual dependencies, spatial dynamics and temporal changes within an image. Three L-3 tasks are included and Figure 6 presents some visualization examples for each task.
@@
  • $\bullet$
    Referring Expression Segmentation: This task involves accurately segmenting and identifying specific regions within an RSI based on a natural language description. It challenges VLMs to translate textual descriptions into pixel-wise segmentation masks.
Attribute Reasoning. In this dimension, VLMs are required to reason about the overall characteristics of RSIs or the specific attributes of the instances. This dimension includes 2 Level-3 tasks, with visualization examples for each task shown in Figure 7.
@@
Assessment Reasoning. In this dimension, VLMs are required to reason about estimated indicators of societal development and environmental conditions depicted in the RSI. This dimension includes 2 Level-3 tasks, with visualization examples for each task shown in Figure 8.
@@
    Resource Assessment: This task evaluates a VLM’s capability to deduce the estimated resource quantity present in a given area, such as the population sizes and economic situations.
@@
Common Sense Reasoning. In this dimension, VLMs are required to reason about real-world scenarios based on the extensive common sense knowledge embedded in LLM. This dimension includes 3 Level-3 tasks, with visualization examples for each task shown in Figure 9.
@@
Table 5: The detailed statistics for CHOICE. Abbreviations adopted: MCQ for Multi-Choice Question; FC for False Color; NL for Nighttime Light; MI for Multi-Image; BT for Bi-Temporal; MT for Multi-Temporal.
  Level-1 Level-2 Level-3 Data Source Answer Type Modality Resolution Image Size # Sample
@@
3/4-option MCQ
RGB
MI
0.3$\sim$30m 512$\times$512 800
MR Natural Earth 4-option MCQ RGB N/A 512$\times$512 157
SC Google Earth Engine 4-option MCQ RGB 0.3$\sim$10m 512$\times$512 2000
IC Google Earth Engine 4-option MCQ RGB 0.3m 512$\times$512 200
SII LR Google Earth Engine 4-option MCQ RGB 0.3m 512$\times$512 100
OC Google Earth Engine 4-option MCQ RGB 0.3m 800$\times$800 500
OL Google Earth Engine 4-option MCQ RGB 0.3m 800$\times$800 500
OP Google Earth Engine 3-option MCQ RGB 0.3m 800$\times$800 500
AR SWISSIMAGE [^4] https://www.swisstopo.admin.ch/de/en/orthoimage-swissimage-10 4-option MCQ RGB 0.1$\sim$0.3m 800$\times$800 500
VG Google Earth Engine Bounding Box RGB 0.3m 512$\times$512 600
HD Google Earth Engine 4-option MCQ RGB 0.3m 800$\times$800 500
CID AC SWISSIMAGE 3-option MCQ RGB 0.1$\sim$0.3m 800$\times$800 350
SR SWISSIMAGE 4-option MCQ RGB 0.1$\sim$0.3m 800$\times$800 300
CD Google Earth Engine 4-option MCQ RGB, BT 0.3m 512$\times$512 200
RES Google Earth Engine Seg. Mask RGB 0.3m 800$\times$800 500
Reasoning AttR TP Sentinel-2 4-option MCQ RGB, MT 10m 512$\times$512 200
@@
Google Earth Engine
CNBH
4-option MCQ RGB 0.1$\sim$0.3m 512$\times$512 300
@@
Google Earth Engine
ODIAC
4-option MCQ RGB 2.4m 512$\times$512 200
@@
3/4-option MCQ
RGB
MI
10m 512$\times$512 600
CSR DD Diffusion Model 4-option MCQ RGB N/A 512$\times$512 200
GD Google Earth Engine 4-option MCQ RGB 0.3m 512$\times$512 200
SI Google Earth Engine 4-option MCQ RGB 0.3$\sim$10m 512$\times$512 300
       Total 10507
A.2 Statistics of CHOICE
All RSIs in CHOICE are sourced from a variety of satellites, products, and platforms. To ensure a balanced evaluation, we targeted an even distribution among problems related to different capabilities during data collection, with a minimum of 150 samples per task. The data sources and the number of problems are listed in Table 5.
Figure 10: Examples of data leakage in DIOR-RSVG dataset. Model abbreviations adopted: Ovis1_6 for Ovis1.6-Gemma2-9B; DeepSeek for DeepSeek-VL-7B; InternVL for InternVL2-8B; LLaVA for LLaVA-1.6-7B.
Appendix B Data Leakage in Current Benchmarks
As shown in Table 1, the majority of mainstream multi-modal datasets and benchmarks in the field of remote sensing are repurposed from commonly used datasets [18, 19], some of which, such as DOTA [20] and DIOR [21], have been utilized during the training stages of VLMs. This practice of data repurposing inevitably leads to data leakage, which compromises the objectivity of the evaluation process. We use DIOR-RSVG [17], which is further processed from the object detection dataset DIOR [21], as an example to demonstrate such data leakage in the context of the visual grounding task. Specifically, one-quarter of the image that contains the object of interest is masked while the query remains unchanged. We then compare the outputs of the VLMs with those obtained from the original images. As illustrated in Figure 10, the results inferred from the masked image still yield an IoU similar to that of the intact images. This reflects the prior knowledge of VLMs on responding even without visual evidence and highlights the inherent data leakage. Therefore, most existing benchmarks derived from such common datasets fail to accurately and objectively reflect the true proficiency of VLMs in remote sensing tasks.
Appendix C Details on CHOICE Construction
Figure 11: Overview of the construction procedure for the Scene Classification task.
To increase data diversity, most RSIs in CHOICE are sourced from 50 randomly selected cities across six continents (excluding Antarctica), based on the top 1,000 global cities identified by Oxford Economics’ Global Cities Index, minimizing potential biases in the data. The details of data coverage are listed in Table 6
@@ Appendix C Details on CHOICE Construction
For Scene Classification, we begin by defining a set of categories for RSIs to guide data collection. OpenStreetMap (OSM) represents physical features on the ground (e.g., roads or buildings) using tags (i.e., key-value pairs), with each tag describing a geographic attribute of the feature. For example, an aerodrome, airport or airfield can be tagged as $aeroway=aerodrome$. Take the common categories in existing relevant datasets [52, 53] as a reference, we follow the OSM’s tagging system to establish a category set, as shown in Table 8, with each predefined category can be described by one or more key-value pairs. We further select 10 of the 50 chosen cities as the primary data collection areas to capture more intra-class variations, including Sapporo, Mecca, Busan, Shanghai, Hamburg, Amsterdam, Phoenix, Cairo, Santiago and Melbourne. As shown in Figure 11, we obtain polygons representing the locations of interest based on the predefined key-value pairs for each category. Using these polygons, we then download the corresponding RSIs from Google Earth Engine. To ensure high-quality data, we design an automatic filtering pipeline to exclude the low-confidence RSIs resulting from the inaccuracy in OSM. Following [14], we finetune two versions of classification models based on the 30-category AID dataset and 45-category NR dataset, both of which include the 20 categories we defined. After cropping the RSIs we collect into $512\times 512$ tiles, we input them into the two models and take the intersection of the outputs with over 99% confidence from both models as the desired RSIs. Finally, human annotators review the results to further ensure the accuracy of all categories. We ultimately collect 2,000 high-confidence RSIs across 20 categories for this task.
For Image Quality, we first define 4 common types of noise that typically interfere with RSIs: Gaussian noise, impulse noise, stripe noise and deadline noise. We randomly select 200 samples in Scene Classification evenly as the original high-quality RSIs, and manually add each type of noise to these samples, creating 50 samples for each noise type, resulting in 200 multiple-choice questions (MCQs) for identifying the type of noise present in an RSI. Moreover, we concatenate each noisy RSI with its corresponding original RSI side by side, forming 200 two-option MCQs to determine which version has better quality. We further collect cloudy RSIs from the Landsat-8 satellite, which has a spatial resolution of 30 meter/pixel, across the 10 cities. By selecting tiles with 60% cloud coverage and cropping the corresponding areas into $512\times 512$ patches, we combine Bands 4, 3 and 2 to obtain the natural color images, which results in 200 cloudy RSIs. We then concatenate three RSIs from Time Property and one cloudy RSI into a $2\times 2$ grid, and generate 200 MCQs for identifying the cloudy RSI. To comprehensively assess the VLM’s capability to identify and compare the quality of RSIs, we concatenate the original RSI, its corresponding noisy version, the cloudy RSI and the low-resolution RSI (obtained by randomly cropping the original RSI and resizing it to the original size) into a $2\times 2$ grid, forming 200 MCQs to identify the best quality among the four options.
For Image Modality, we include four modalities of RSIs: RGB, SAR, false color and nighttime light. We still select the RSIs from Scene classification as the RGB group, and combine Bands 5, 4 and 3 of the cloudy RSIs from Landsat-8 as the false color modality. With the assistance of the Sentinel-1 satellite with a spatial resolution of 10 meter/pixel, we download and crop SAR RSIs (both VV and VH polarizations) across the 10 cities via AI Earth [^6] https://engine-aiearth.aliyun.com. As for nighttime light images, we manually choose RSIs from the SDGSAT-1 satellite. We then randomly select 200 RSIs per modality to form 800 MCQs for identifying the modality of each RSI.
For Image Caption, we randomly select 200 RSIs from Scene Classification ensuring an even distribution, and employ human annotators to generate a descriptive caption for each RSI. Then we prompt GPT-4 to generate 3 distracting descriptions, with these descriptions subsequently verified by human reviewers to ensure accuracy and relevance. The prompt we use is shown in Figure 12.
For Landmark Recognition, we choose 100 most famous landmarks in the world from List Challenges [^7] https://www.listchallenges.com/ and download the RSIs with 0.3 meter/pixel spatial resolution from Google Earth Engine according to their latitude and longitude coordinates.
For the other 6 leaf tasks in Single-instance Identification dimension, objection detection for identifying instances and obtaining their coordinates of bounding boxes is the first step to generate associated problems. As mentioned in Scene Classification, we also predefine a set of categories for instances based on the existing object detection datasets [20, 21] and OSM key-value pairs, as shown in Table 9. We select 30 of the 50 chosen cities as the primary data collection areas to capture more intra-class variations, including Riyadh, Gwangju, Tokyo, Daejeon, Chengdu, Istanbul, Nagoya, Wuhan, Hsinchu, Nanjing, Kuala Lumpur, Jeddah, Lyon, Paris, Rotterdam, Brussels, Berlin, Munich, Vienna, Zurich, Atlanta, Boston, Washington D.C., Montreal, Calgary, Windhoek, Gaborone, Rabat, Montevideo and Lima. Using the same procedure as for Scene Classification, we obtain the original RSIs from Google Earth Engine and crop them into $800\times 800$ patches. We apply two object detection models, MTP [37] pretrained on DIOR-R [54] and YOLOv8 [38] pretrained on DOTA [20], to automatically generate bounding boxes for the predefined categories. Human annotators are then employed to verify the completeness and accuracy of these labels, retaining only those with high confidence from both models. Once getting the labels and coordinates of these instances, we count the number of specific objects for Object Presence and Object Counting. We compute the centroids of these objects and assign their coarse locations in a $3\times 3$ grid for Object Localization, and use the normalized coordinates of bounding boxes directly as the label for Visual Grounding. For Hallucination Detection, we aim to assess the VLM’s capability to realize the non-existent instances within RSIs and correctly refuse to respond to such queries.
For Attribute Recognition, RSIs collected from Google Earth Engine with a resolution of 0.3 meter/pixel are not sufficient for fine-grained attributes such as vehicle color. Therefore, we collect additional RSIs with a higher resolution of 0.1 meter/pixel. We then utilize K-means clustering to classify the color of the vehicles and obtain the direction by the angle of bounding box, generating MCQs related to these attribute.
For Change Detection, we collect registered RSIs captured at two different timestamps from 6 of the 50 cities, including Nicosia, Haifa, Dammam, Malmo, Seattle and Pretoria. Human annotators are enlisted to generate a descriptive caption for each pair of RSIs that highlights the temporal changes between them. GPT-4 is then prompted to generate 3 distracting options, with these descriptions subsequently verified by human reviewers to ensure accuracy and relevance. The prompt we use is shown in Figure 12.
For Referring Expression Segmentation, we follow the method mentioned in Single-instance Identification. Instead of applying object detection foundation models, we employ a segmentation model, RingMo-SAM [39], to generate precise segmentation masks for the objects of interest, with the bounding boxes serving as the visual prompts to guide the segmentation process, ensuring more accurate delineation of the objects. By computing and comparing the locations or areas of these object masks, we are able to automatically construct RES problems, such as "Give me the segmentation mask of the larger roundabout."
For Physical Property, we utilize the CNBH-10m [56] product, which provides the estimated building heights in China, to guide the collection of RSIs. We randomly select 100 points from this product and download the corresponding RSIs from Google Earth Engine. To ensure accuracy, volunteers compare the estimated building heights provided by CNBH-10m with human-assessed heights from the RSIs. Any discrepancies are corrected by human annotators to maintain accuracy in building height measurements. Moreover, we incorporate an additional 200 MCQs to determine whether vehicles are stationary or in motion, based on the images collected in Attribute Recognition.
Appendix D Details on Evaluation
Table 7: Details of the evaluated open-source VLMs.
  Model Vision Backbone Language Backbone Overall Parameters
       General-domain Large Vision-Language Models
Qwen2-VL-7B ViT-L Qwen2-7B 7B
Qwen2-VL-72B ViT-L Qwen2-72B 72B
Ovis1.6-Gemma2-9B Siglip-400M Gemma2-9B-It 9B
InternVL2-8B InternViT-300M-448px internlm2_5-7b-chat 8.1B
InternVL2-26B InternViT-6B-448px-V1-5 internlm2-chat-20b 25.5B
InternVL2-40B InternViT-6B-448px-V1-5 Nous-Hermes-2-Yi-34B 40.1B
LLaVA-1.6-7B CLIP ViT-L/14 Vicuna-v1.5-7B 7.1B
LLaVA-1.6-13B CLIP ViT-L/14 Vicuna-v1.5-13B 13.4B
Llama3.2-11B - Llama 3.1 10.6B
GLM-4V-9B - - 9B
DeepSeek-VL-7B SAM-B&SigLIP-L DeekSeek-7B 7.3B
MiniCPM-V-2.5 Siglip-400M Llama3-8B-Instruct 8.5B
Phi3-Vision CLIP ViT-L/14 Phi-3-mini 4.2B
mPLUG-Owl3-7B Siglip-400M Qwen2-7B 8B
Molmo-7B-D CLIP ViT-L/14-336px Qwen2-7B 7B
Remote Sensing Large Vision-Language Models
GeoChat CLIP ViT-L/14-336px Vicuna-v1.5-7B 7B
LHRS-Bot CLIP ViT-L/14 LLaMA2-7B 7B
LHRS-Bot-nova SigLIP-L/14-336px LLaMA3-8B 8B
VHM CLIP ViT-L/14-336px Vicuna-v1.5-13B 7B
Table 7 provides details of all open-source VLMs evaluated in CHOICE.
By leveraging the instruction-following ability of VLMs, we restrict the model response to either the A/B/C/D option or the corresponding content of the choice. This constraint simplifies subsequent choice extraction and enables straightforward accuracy computation. As shown in Figure 13, we design a system prompt tailored for both general-domain VLMs and RSVLMs when handling tasks with MCQs. For Visual Grounding, where the response requires free-form coordinates, we retain the original system prompts specific to RSVLMs, as these are better suited for generating precise coordinate outputs. This dual approach ensures optimal performance across varied task requirements while maintaining consistency and accuracy in evaluation.
We also present an example of evaluation strategy for CLIP-based VLMs in Figure 13. The MCQ is transformed into $n$ declarative sentences by adding or replacing keywords in the question and further rearranging the word order. During the evaluation stage, the similarity score between the RSI and each sentence is computed. The choice corresponding to the sentence that achieves the highest similarity score is considered the model’s response. Notably, CLIP-based VLMs lack the capability for visual grounding, and directly computing similarity scores between RSIs and pure coordinates is not feasible. Consequently, this leaf task is left unaddressed for CLIP-based VLMs.
Failure Cases. For Visual Grounding, the responses from LHRS-Bot consistently fall within a limited set of repeated outputs across all problems, such as $(0.44,0.44,0.52,0.52)$, $(0.34,0.44,0.66,0.66)$ and $(0.38,0.38,0.59,0.59)$. This pattern results in extremely low accuracy and highlights a significant lack of diversity and adaptability in the model’s predictions for this task. For Attribute Comparison, LHRS-Bot consistently outputs the choice “No”, regardless of how the choice order is altered. This behavior indicates a fundamental lack of capability in the model to perform attribute comparison tasks effectively.